OpenStudy (anonymous):

write the equation in slope-point form for a line that contains the points (1,-3) and (-5, 2) Please show your steps

OpenStudy (anonymous):

@ConfusedAboutCalculus I thought you confused about Calculus. This is basic algebra 1! You should know the slope formula of linear equation?

hero (hero):

Hint use y = mx + b

hero (hero):

Insert both points into y = mx + b to create two separate equations: -3 = 1m + b 2 = -5m + b

hero (hero):

Now you have a system of equations. Multiply the first equation by 5 to get -15 = 5m + b 2 = -5m + b

hero (hero):

Combine both equations together to get -13 = 2b

hero (hero):

Divide both sides by 2 to get: \(-\large\frac{13}{2} = b\)

hero (hero):

Now substitute the value for b back into one of the original equations: \(-3 = 1m - \frac{13}{2}\)

hero (hero):

Now add 13/2 to both sides: \(\frac{13}{2} - 3 = m\) Now simplify to solve for m: \(\frac{13}{2} - \frac{6}{2} = m\) \(\frac{7}{2} = m\)

hero (hero):

Now that you've found m and b, rewrite the general form y = mx + b, but include the actual values of m and b that you found: \(y = \frac{7}{2}x - \frac{13}{2}\)

hero (hero):

Me personally, I like to combine the right side into one fraction to make it even simpler: \(y = \frac{7x - 13}{2}\)